# When do World Models Successfully Learn Dynamical Systems?

### v1

@@ 4 Datasets and numerical settings
In order to verify the theoretical results of Section 3, we created datasets for the heat, wave, and KS equations. To verify the methodology on a standard computational fluid dynamics problem, we also included a published dataset for flow around a cylinder at Reynolds number 3900, a well-studied case in the literature. The datasets are learned with a variety of methods, starting with a simple least squares regression model, to gradient descent methods, through to full DCGAN methods employed by NVIDIA’s LongVideoGAN [48].
It is important to point out that all of the datasets used in this work are fundamentally dynamical, i.e. they are videos. We give some static snapshots (Figs. 9, 10, 12, 16) but in order to get a proper impression of the both the dataset and model outputs, the reader is strongly encouraged to examine the videos in the supplementary material.
The heat, wave and KS datasets use a Gaussian random field as the initial condition of the flow field $u$. Where applicable, the initial condition of the velocity field $v$ was set to be identically 00. All three datasets use periodic boundary conditions. The precise numerical parameters used to generate the Gaussian random field are listed as Table 1 in Appendix A.
Datasets for linear equations.
The heat and wave equation solutions are constructed by forward Euler, an explicit method. This scheme, while simple, is known to converge for appropriately small time steps for these equations [49]. Since the Laplacian $\nabla^{2}$ is not observable under the tokenization operator (Theorem 6) we generate a modified Laplacian $\nabla(a\nabla)$ where $a$ is a stochastically generated positive definite diagonal matrix. Examples of the low-res (super-res) heat (wave) dataset can be found in the left hand column of Fig. 10 (Fig. 12), in the Appendix. The compression factor was selected to roughly match that of the LongVideoGAN model. Table 2 gives further details on the production of these datasets.
Kuramoto-Sivashinsky Equation.
A large amount of data was created by adopting the ETDRK4 (an exponential 4th order Runge-Kutta method) [50], which we modified to produce solutions to the 2D equation. Notably, the 0th mode in Fourier space was zeroed out to prevent the growth of the mean $\bar{u}$ (see Eq. 2.2 in [51]). The precise numerical parameters used can be found in Table 3, and still frames from the super-res dataset can be found in the left column in Fig. 9.
Flow Around a Cylinder.
For the experiments involving flow around a cylinder at Reynolds number 3900, we used a publicly available dataset [52] provided by the authors of [10]. The dataset consists of 100,000 greyscale images with a resolution of $1000\times 600$ pixels, capturing the characteristic Kármán vortex street - a coherent structure of alternating vortices aligned with the cylinder’s axis. The data was generated using Large Eddy Simulation (LES), and the resulting unsteady velocity fields were processed via a projection mapping technique that preserves ergodicity [53] in a reduced state space. For details on the LES numerical setup, we refer to [10]. The details on the setup and configuration of neural network training can be found in Section A.7.
@@ 5.1 Linear Equations
Here we present the results of the training on the data produced for the heat and wave equations, as in Eq. 11. We learn the operator both by using modern gradient descent learning methods and traditional multivariate linear regression, to verify the machine learning methods. From Thm. 7 we know the reconstruction map exists, we just have to ensure the model can in fact find it in practice.
5.1.1 Least-Squares Regression
We begin in a fully specified statistical model for the low resolution model, where the learning is done by simple multi-variate least squares regression. Fig.15 shows the average absolute error (of the residues) when the history length of the model is varied, whilst Fig. 15 shows the maximum error across all pixels in all 2000 frames. The model is trained on a dataset of 100 initial conditions, and evaluated on a new, unseen initial condition. Each point is sampled 100 times. The shaded area represents one standard deviation from the mean. Note the significant drop in error; we are guaranteed by Lemma 2 that this occurs as the supplied history length to the model exceeds the compression factor $k$. We can see that in this ‘ideal’ case (optimality given by Gauss-Markov theorem) the drop actually happens much earlier.
5.1.2 Autoregressive Learning
Fig. 2 shows how the error evolves against steps for stochastic gradient descent via the Adam [54] optimiser. The model shows a clear improvement as history size increases, but past the compression factor of $16$, increasing the history size provides no benefit, as discussed in Section 3. In particular, the lines representing a history size of $16$ and $20$ have converged by the end of training. Note also that the model performs poorly for an unmodified (non-observable) heat operator.
5.1.3 Full Pipeline Learning
@@ 5.2 Non-linear equations
Loss curves for frame-wise $L_{1}$ and $L_{2}$ residue loss can be found in Fig. 5. We computed $\rho(\Delta t)$ (See Section A.7.3) for $15$ generated and $15$ real $300$ frame videos from the test set. We compare the sample means and standard deviation in Fig. 5. Note the model has learned this statistic well for early times, but appears to ‘desync’ at longer time horizons.
Figure 2: Average $L_{2}$ residues of our low-res heat equation model against epoch, for the test set. Each x-axis unit represents 2000 epochs. Data is normalised between $[-1,1]$. Here, the unmodified (constant coefficient $a$, non-observable) heat equation has a history size of $16$.
@@
Figure 4: (Left) Residues of our KSE model against epoch, for the low-res test set. For data normalised between $[-1,1]$, We obtain a minimum average $L_{2}$ loss of $4\times 10^{-6}$ and minimum average $L_{1}$ loss of $0.0014$.
Figure 5: (Right) Temporal correlations for the point (128, 128) for the trained model and ground truth. The shaded area represents 1 standard deviation from the mean. The statistics are calculated across 15 real and 15 generated videos. The videos were produced as in Fig. 3; 16 frames are given to the low-res model, a full 300 frames are produced, and the results handed over to the super-res model.
Figure 6: $\Delta t$ vs $\rho(\Delta t)$ with $\Delta t\in{1,\dots,100}$ for pixel 1 (left), and pixel 3 (right). The lines represent the mean correlations and the shaded areas show a 1 standard deviation range. The mean and standard deviation of the correlation are for the generated videos are calculated across $50$ generated videos containing $301$ frames each, while the statistics for the real videos were calculated across $100$ real videos with $301$ frames.
5.3 Flow around a cylinder
@@
We computed $\rho(\Delta t)$ for $\Delta t\in\{1,\dots,100\}$ and for $4$ different choices of $P$ for $100$ non-overlapping videos with a length of $301$ frames of the original data, as well as for $50$ different generated videos with a length of $301$ frames. We then computed the sample means and standard deviations of the correlation coefficients to compare our generated videos to the original ones. Pixels $1$, $2$, $3$ and $4$ used for comparison are given in Fig. 16.
@@
The results in Fig. 6 indicate that, overall, the generated videos exhibit temporal dynamics consistent with the real videos, as the mean correlations almost always fall within one standard deviation of the real data. However, the model tends to exaggerate short-term correlations, particularly for the first Pixels $1$, $2$ and $3$, showing higher magnitudes of local extrema for small values of $\Delta t$. Long-term correlations are captured effectively for the most part, but there are notable deviations at Pixel $4$, where the generated videos diverge significantly from the real videos for $\Delta t>50$, indicating difficulty in replicating more complex, less periodic behaviours that tend to happen on the right-hand side of the image. An analysis of all four coordinate points and longer, $1000$ frame videos is given in Section A.10.
5.4 Computational Cost
@@
The motivation to generate dynamical systems using GANs stems in part from their ability to significantly reduce computational costs at inference time compared to LES. Exact training times and hardware are listed for each model in Section A.8.
@@ Conclusion
We have introduced a system-theoretic framework for understanding when ‘world models’ can faithfully learn dynamics from tokenized observations. By relating observability to the existence of an autoregressive map in latent space, we explained when concatenated token histories suffice for reconstruction in both linear and nonlinear settings. Empirically, our sequence of models from analytic least-squares to full GANs validated this theory across four benchmark problems: modified heat and wave equations, the chaotic 2D Kuramoto–Sivashinsky equation, and a high-Reynolds CFD dataset of a Kármán vortex street. In each case, world-model reconstructions achieved low prediction error and captured key temporal correlations. Compared to traditional LES, inference with our generative models is orders of magnitude faster.
Limitations
@@
Our analysis relies on a strong global observability assumption: the output map and underlying flow must admit a continuous, invertible reconstruction operator. In practice, certain coefficient choices (e.g., constant conductivity in the heat equation) violate observability under patch averaging, leading to degraded performance unless randomness is injected. Moreover, while world models generalize well within the distribution of training trajectories, their accuracy degrades over very long rollouts. In our work, we see this occur for the chaotic regimes in terms of residues, but this is not necessarily an issue, since small variations in initial conditions lead to large variations later in time, a problem even the LES suffers from, while residues can only measure local reconstruction error. However, we still see a breakdown in the time correlation for longer roll outs, especially in the KSE case. Finally, since observability is a binary conditions, further research is required to characterise when exactly a dataset may satisfy the formal assumptions set out in Section 3, but in practice still be difficult to learn i.e. in some sense be ‘close’ to being non-observable.
Acknowledgments and Disclosure of Funding
@@
This work was funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within the priority program "Carnot Batteries: Inverse Design from Markets to Molecules" (SPP 2403) grant no. 526152410 and in another part by the Math+ Center of Excellence as an activity in the project "AI Based Simulation of Transient Physical Systems – From Benchmarks to Hybrid Solutions". We also acknowledge support from the NVIDIA Academic Grant Program in the form of compute resources. The authors thank Onur T. Doganay, Martin Eigel, Yuan Huang and Alexander Klawonn for interesting discussions.
Appendix A Appendix: Supplementary Material to ‘When do World Models Successfully Learn Dynamic Systems’?
A.1 Extraneous Proofs
@@
Here we give detail on some proofs of Section 3. We first formally define the empirical risk functions defined in Section 3.2. The empirical risk functions are defined by
^$\hat{\mathcal{L}}_{g,N,T}(\eta_{g})$ $=\frac{1}{N(\frac{T}{\Delta}-k-1)}\sum_{j=1}^{N}\sum_{t\in\Gamma_{T}\setminus(\Gamma_{\Delta\bar{k}}\cup\{T\})}\|\eta_{g}(y_{\mathrm{seq}}(t;x_{0}^{(j)}))-y(t+\Delta;x_{0}^{(j)})\|^{2},$ (15a)
^$\hat{\mathcal{L}}_{G,N,T}(\eta_{G})$ $=\frac{1}{N(\frac{T}{\Delta}-k)}\sum_{j=1}^{N}\sum_{t\in\Gamma_{T}\setminus\Gamma_{\Delta k}}\|\eta_{G}(y_{\mathrm{seq}}(t;x_{0}^{j}))-x(t;x_{0}^{(j)})\|^{2}.$ (15b)
@@
Here $T$ stands for the observation time of the trajectory, $\Delta$ for the time steps of the evolution and observation, and $N$ is the number of initial conditions $x_{0}^{(i)}$ sampled from some distribution $\mu$ on $\mathcal{X}$. Clearly,
$\mathbb{E}_{x_{0}\sim\mu}[\hat{\mathcal{L}}_{N,T,g}(\eta_{g})]=\mathcal{L}_{T,g}(\eta_{g})\text{ and }\mathbb{E}_{x_{0}\sim\mu}[\hat{\mathcal{L}}_{T,G}(\eta_{G})]=\mathcal{L}_{T,G}(\eta_{G}).$
@@
We give the proof for the autoregressive map $g:\mathcal{Y}^{k}\to\mathcal{Y}$, the proof of PAC-learnability for $G:\mathcal{Y}^{k}\to\mathcal{X}$ is completely analogous.
@@
00 0 $\leq\mathcal{L}_{g}(\hat{\eta}_{g,N,T})=\mathcal{L}_{g}(\eta_{g})+\left[\hat{\mathcal{L}}_{g,N,T}(\hat{\eta}_{g,N,T})-\hat{\mathcal{L}}_{g,N,T}(\eta_{g})\right]$
@@
That this hypothesis space exists follows from the universal approximation property of neural networks, see e.g. [55, 56, 57] which guarantee $\|\eta_{g}-g\|_{\infty}\leq\sqrt{\frac{\varepsilon}{2}}$ for some $\eta_{g}\in\mathcal{H}_{\varepsilon}$. From this and by (7a).
@@
Not that all summands occurring in ^$\hat{\mathcal{L}}_{g,N,T}(\eta_{g})$ are uniformly bounded for $\eta\in\mathcal{H}_{\varepsilon}$ and continuous in the neural networks’ weights $\theta\in\Theta_{\varepsilon}$. By the uniform law of large numbers [58] it follows that the second term in (16) converges to zero $P$-almost surely when $\mathcal{H}_{N,T}$ is replaced by $\mathcal{H}_{\varepsilon}$. As almost sure convergence implies convergence in probability, there exists a $N(\varepsilon,\delta)$ sufficiently large such that for $N\geq N(\varepsilon,\delta)$
@@
Here we have given the proof of Theorem 4 in the large sample limit $N\to\infty$ for i.i.d. sampling of the initial conditions $x^{(j)}_{0}\sim\mu$. For ergodic dynamic systems, by application of Birkhoff’s ergodic theorem [59], this limit can be replaced by the large time limit $T\to\infty$ and only one observed trajectory $N=1$, when $x_{0}\sim\mu$ and $\mu$ is the invariant measure of the dynamic system, see [10] for a similar argument.
@@
$\begin{split}&P\left(\mathcal{L}_{G,T}(\hat{\eta}_{G,N,T})>\frac{\varepsilon}{6\left\|\frac{d(S^{-k}_{\Delta})_{*}\mu}{d\mu}\right\|_{\infty}\left(\frac{T}{\Delta}-k\right)}\right)\leq\frac{\delta}{2}\\ &\mbox{ for }N\geq N_{G}\left(\frac{\varepsilon}{6\left\|\frac{d(S^{-k}_{\Delta})_{*}\mu}{d\mu}\right\|_{\infty}\left(\frac{T}{\Delta}-k\right)},\frac{\delta}{2}\right).\end{split}$ (21)
@@
$\begin{split}&P\left(\mathcal{L}_{g,T}(\hat{\eta}_{g,N,T})>\frac{\varepsilon}{6\left\|\frac{d(S^{-k}_{\Delta})_{*}\mu}{d\mu}\right\|_{\infty}\left(\frac{T}{\Delta}-k\right)}\right)\leq\frac{\delta}{2}\\ &\mbox{ for }N\geq N_{g}\left(\frac{\varepsilon}{6L_{\mathcal{H}_{G,N,T}}\left\|\frac{d(S^{-k}_{\Delta})_{*}\mu}{d\mu}\right\|_{\infty}\left(\frac{T}{\Delta}-k\right)},\frac{\delta}{2}\right).\end{split}$ (22)
Using along with (LABEL:eq:LipshitzEta_G) and (LABEL:eq:Est_g), we also see that the second term on the right hand side of (LABEL:eq:ErrorDecompFullState) is larger $\frac{\varepsilon}{2}$ with probability at most $\frac{\delta}{2}$ and for $N\geq N_{g}$. Therefore, for $N\geq\max\{N_{g},N_{G}\}$, we obtain the assertion of Theorem 5 from (LABEL:eq:ErrorDecompFullState), the union bound for probabilities and equations (LABEL:eq:ProbBound_g) and (LABEL:eq:ProbBound_g).
@@
Let $\Gamma$ be the quadratic lattice with lattice constant $\delta>0$, edge length $L$ and $\sqrt{k}=\frac{L}{\delta}$ lattice points on the edge. Eigenvectors of $A_{h}$ for $a=\text{const.}$. are those of the standard lattice Laplacian, which are well known. In fact, the trigonometric functions $(\sin(\omega\cdot\xi))_{\xi\in\Gamma}$, $(\cos(\omega\cdot\xi))_{\xi\in\Gamma}$ diagonalise $A_{h}$ where $\omega=(\omega_{1},\omega_{2})$ and $\omega_{j}\in\{2\frac{\pi j\delta}{L}:j\in\{0,\ldots,\sqrt{k}-1\}\}$.

### v2

@@ 4 Datasets and numerical settings
In order to verify the theoretical results of Section 3, we created datasets for the heat, wave, and KS equations. To verify the methodology on a standard computational fluid dynamics problem, we also included a published dataset for flow around a cylinder at Reynolds number 3900, a well-studied case in the literature. More details are given in Section G. The datasets are learned with a variety of methods, starting with a simple least squares regression model, to gradient descent methods, through to full DCGAN methods employed by NVIDIA’s LongVideoGAN (Long_video_gan).
It is important to point out that all of the datasets used in this work are fundamentally dynamical, i.e. they are videos. We give some static snapshots (Figs. 9, 10, 11, 15) but in order to get a proper impression of the both the dataset and model outputs, the reader is strongly encouraged to examine the videos in the supplementary material.
The heat, wave and KS datasets use a Gaussian random field as the initial condition of the flow field $u$. Where applicable, the initial velocity field $v$ was identically set to 0. All three datasets use periodic boundary conditions. The precise numerical parameters used to generate the Gaussian random field are listed as Table 6 in Appendix G.
@@ 5.1 Linear Equations
Figure 2: Average $L_{2}$ residues of our low-res heat equation model against epoch, for the test set. Each x-axis unit represents 2000 epochs. Data is normalised between [1,1][-1,1]. The unmodified (non-observable) heat equation has history size $16$.
Here we present the results of the training on the data produced for the heat and wave equations, as in Eq. 10. We learn the operator both by using modern gradient descent learning methods and traditional multivariate linear regression, to verify the machine learning methods. Thm. 7 guarantees the reconstruction map exists; it remains to check it can be found in practice.
Least-Squares Regression.
We begin in a fully specified statistical model for the low resolution model, where the learning is done by multivariate least squares regression. Fig.14 shows the average absolute error of the residues when the history length of the model is varied, whilst Fig. 14 shows the maximum error across all pixels in all 2000 frames. Each point is sampled 100 times. The shaded area represents a $1\sigma$ range. Note the significant drop in error; we are guaranteed by Lemma 2 that this occurs as the supplied history length to the model exceeds the compression factor $k$. We can see that in this ‘ideal’ case (optimality given by Gauss-Markov theorem) the drop actually happens much earlier.
Autoregressive Learning.
Fig. 2 shows how the error evolves against steps for stochastic gradient descent via the Adam (adam) optimiser. The model shows a clear improvement as history size increases, but past the compression factor of $16$, increasing the history size provides no benefit, as discussed in Section 3. In particular, the lines representing a history size of $16$ and $20$ have converged by the end of training. Note also that the model performs poorly for an unmodified (non-observable) heat operator. This demonstrates that observability is a necessary condition for learning, as stated in Section 3, and adding more information beyond simply satisfying observability doesn’t change the end result.
Full Pipeline Error.
@@ 5.2 Non-linear equations
Loss curves for frame-wise $L_{1}$ and $L_{2}$ residue loss can be found in Fig. 5. We computed $\rho(\Delta t)$ (See Section H.3) for $15$ generated and $15$ real $300$ frame videos from the test set. We compare the sample means and standard deviation in Fig. 5. Note the model has learned this statistic well for early times, but appears to ‘desync’ at longer time horizons.
@@
Figure 4: (Left) Residues of our KSE model against epoch, for the low-res test set. For data normalised between [1,1][-1,1], We obtain a minimum average $L_{2}$ loss of $4\times 10^{-6}$ and minimum average $L_{1}$ loss of $0.0014$.
Figure 5: (Right) Temporal correlations at the point (128, 128) for the trained model and ground truth. The shaded areas show a $1\sigma$ range. 15 videos were produced as in Fig. 3; the low-res and super-res models are concatenated.
Comparison to Existing Models.
A direct comparison of our method with li2020fourier and lu2021learning can be found in Table 1. Both were trained in a data-driven manner; the DeepONet uses an FNO as the branch network and an FCN as the trunk. For the linear equations, we compute the MSE and L1 (error) norm for a 1, 200, and 1000 frame look-ahead prediction. Initial conditions were sampled random fields distinct from the training data, but with the same distribution (in-distribution). For out-of-distribution initial conditions, see Table 3 in the Appendix. The KSE also includes a 1 frame look-ahead MSE and L1 norm, as well as a metric measuring the distance between the correograms (x-axis 1-200 frames) of the model and ground truth data (Table 2). Here, ‘diverged’ refers to a value larger than float precision. We note that our model obtains better performance metrics in the majority of cases. This is especially true of long time horizons, which underlines the stabilising effect of our history based autoregressive approach. Further comparisons can be found in Section C.
We omit a comparison to PINNs (raissi2019physics), as PINNs need to be retrained for new initial conditions, which affects robustness and speed at generation time. Additionally, the KVS dataset contains partial information of the state, and thus the data driven case is more general. Current experimental methods (e.g. as in Ghazijahani_2024) also only give a partial measurement. The theoretical guarantees given by Theorems 4 and 5 make sure that despite this, assuming observability, the physics is learned accurately.
5.3 Flow around a cylinder.
@@
We computed $\rho(\Delta t)$ for $\Delta t\in\{1,\dots,100\}$ and for $4$ different choices of $P$ for $100$ non-overlapping videos with a length of $301$ frames of the original data, as well as for $50$ different generated videos with a length of $301$ frames. We then computed the sample means and standard deviations of the correlation coefficients to compare our generated videos to the original ones. Pixels $1$, 22, 33 and $4$ used for comparison are given in Fig. 15.
@@
The results in Fig. 6 indicate that, overall, the generated videos exhibit temporal dynamics consistent with the real videos, as the mean correlations almost always fall within one standard deviation of the real data. However, the model tends to exaggerate short-term correlations, particularly for points $1$, 22 and $3$. Whilst long-term correlations are captured effectively, there are notable deviations at late times at point $4$, suggesting difficulty in replicating more non-periodic behaviours. An analysis of all four coordinate points and longer, $1000$ frame videos is given in Section K.
Figure 6: $\Delta t$ vs $\rho(\Delta t)$ with $\Delta t\in{1,\dots,100}$ for pixel 1 (left), and pixel 3 (right). The shaded areas show a $1\sigma$ range. The mean and standard deviation of the correlation are calculated across $50$ generated videos containing $301$ frames, and for the real videos were across $100$ real videos with $301$ frames.
Ours (full model) DeepONet FNO
L1 MSE L1 MSE L1 MSE
Heat (1) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.0370}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.0021}$} $0.2754$ $0.1181$ $0.2753$ $0.1180$
Heat (200) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.0243}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.0009}$} $0.3204$ $0.1603$ $0.2546$ $0.0996$
Heat (1000) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.0051}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{4.27\times 10^{-5}}$} Diverged Diverged Diverged Diverged
Wave (1) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.0921}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.0133}$} $0.2587$ $0.1090$ $0.2588$ $0.1090$
Wave (200) $0.3181$ $0.1600$ $5.1761$ $2787.4028$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.2741}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.1186}$}
Wave (1000) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.4676}$} $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.3420}$} Diverged Diverged Diverged Diverged
KSE (1) $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.0095}$} $0.0005$ $0.0259$ $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbf{0.0005}$} $0.0190$ $0.0094$
Table 1: Comparison of our model’s performance to state-of-the-art neural operator methods across our datasets. Performance after 1, 200 and 1000 frames of continuous generation is given. Best results for each metric are highlighted in bold.
Computational Cost.
@@
The motivation to generate dynamical systems using GANs stems in part from their ability to significantly reduce computational costs at inference time compared to LES. Exact training times and hardware are listed for each model in Section I.
@@ Conclusion.
We have introduced a system‑theoretic operator learning framework for understanding when ‘world models’ can faithfully learn dynamics from tokenized observations. With observability, we explained when concatenated token histories suffice for reconstruction in both linear and nonlinear settings. Empirically, our sequence of models from analytic least‐squares to full GANs validated this theory across four benchmark problems. In each case, world‑model reconstructions achieved low prediction error and captured key temporal correlations. Compared to traditional LES, inference with our generative models is orders of magnitude faster. Compared to existing neural operator methods, our models perform much better for longer time horizons.
Limitations.
@@
Our analysis relies on a strong global observability assumption: the output map and underlying flow must admit a continuous, invertible reconstruction operator. In practice, certain coefficient choices violate observability under patch averaging. Moreover, accuracy of World Models can degrade over very long rollouts. We see this occur for the chaotic regimes, but this is not necessarily an issue since small variations in initial conditions lead to large variations later in time, a problem even the LES (and neural operators) suffer from. However, we still see a breakdown in the time correlation for longer roll outs, especially in the KSE case. Finally, since observability is a binary conditions, further research is required to characterise when exactly a dataset may satisfy the formal assumptions set out in Section 3, but in practice still be difficult to learn i.e. in some sense be ‘close’ to being non-observable.
6.0.1 Acknowledgements
@@
This work was funded in part by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) within the priority program "Carnot Batteries: Inverse Design from Markets to Molecules" (SPP 2403) grant no. 526152410 and in part by the Math+ Center of Excellence as an activity in the project "AI Based Simulation of Transient Physical Systems – From Benchmarks to Hybrid Solutions". We also acknowledge support from the NVIDIA Academic Grant Program in the form of compute resources. The authors thank Onur T. Doganay, Martin Eigel, Yuan Huang and Alexander Klawonn for interesting discussions.
Appendix A Extraneous Proofs
@@
Here we give detail on some proofs of Section 3. We first formally define the empirical risk functions defined in Section 3.2:
$\hat{\mathcal{L}}_{g,N,T}(\eta_{g})$ $=\frac{1}{N(\frac{T}{\Delta}-k-1)}\sum_{j=1}^{N}\sum_{t\in\Gamma_{T}\setminus(\Gamma_{\Delta\bar{k}}\cup\{T\})}\|\eta_{g}(y_{\mathrm{seq}}(t;x_{0}^{(j)}))-y(t+\Delta;x_{0}^{(j)})\|^{2},$
$\hat{\mathcal{L}}_{G,N,T}(\eta_{G})$ $=\frac{1}{N(\frac{T}{\Delta}-k)}\sum_{j=1}^{N}\sum_{t\in\Gamma_{T}\setminus\Gamma_{\Delta k}}\|\eta_{G}(y_{\mathrm{seq}}(t;x_{0}^{j}))-x(t;x_{0}^{(j)})\|^{2}.$
@@
Here $T$ stands for the observation time of the trajectory, $\Delta$ for the time steps of the evolution and observation, and $N$ is the number of initial conditions $x_{0}^{(i)}$ sampled from some distribution $\mu$ on $\mathcal{X}$. They take their minimums when
Clearly,
[…]
@@
We give the proof for the autoregressive map $g:\mathcal{Y}^{k}\to\mathcal{Y}$; the proof of PAC-learnability for $G:\mathcal{Y}^{k}\to\mathcal{X}$ is completely analogous.
@@
$0$ $\leq\mathcal{L}_{g}(\hat{\eta}_{g,N,T})=\mathcal{L}_{g}(\eta_{g})+\left[\hat{\mathcal{L}}_{g,N,T}(\hat{\eta}_{g,N,T})-\hat{\mathcal{L}}_{g,N,T}(\eta_{g})\right]$
@@
That this hypothesis space exists follows from the universal approximation property of neural networks, see e.g. cybenko1989approximation; yarotsky2017error; kidger2020universal which guarantee $\|\eta_{g}-g\|_{\infty}\leq\sqrt{\frac{\varepsilon}{2}}$ for some $\eta_{g}\in\mathcal{H}_{\varepsilon}$. From this and by (7a).
@@
Not that all summands occurring in $\hat{\mathcal{L}}_{g,N,T}(\eta_{g})$ are uniformly bounded for $\eta\in\mathcal{H}_{\varepsilon}$ and continuous in the neural networks’ weights $\theta\in\Theta_{\varepsilon}$. By the uniform law of large numbers (ferguson2017course) it follows that the second term in (15) converges to zero $P$-almost surely when $\mathcal{H}_{N,T}$ is replaced by $\mathcal{H}_{\varepsilon}$. As almost sure convergence implies convergence in probability, there exists a $N(\varepsilon,\delta)$ sufficiently large such that for $N\geq N(\varepsilon,\delta)$
@@
Here we have given the proof of Theorem 4 in the large sample limit $N\to\infty$ for i.i.d. sampling of the initial conditions $x^{(j)}_{0}\sim\mu$. For ergodic dynamic systems, by application of Birkhoff_proof’s ergodic theorem, this limit can be replaced by the large time limit $T\to\infty$ and only one observed trajectory $N=1$, when $x_{0}\sim\mu$ and $\mu$ is the invariant measure of the dynamic system, see drygala2022generative for a similar argument.
@@
$\begin{split}&P\left(\mathcal{L}_{G,T}(\hat{\eta}_{G,N,T})>\frac{\varepsilon}{6\left\|\frac{d(S^{-k}_{\Delta})_{*}\mu}{d\mu}\right\|_{\infty}\left(\frac{T}{\Delta}-k\right)}\right)\leq\frac{\delta}{2}\\ &\mbox{ for }N\geq N_{G}\left(\frac{\varepsilon}{6\left\|\frac{d(S^{-k}_{\Delta})_{*}\mu}{d\mu}\right\|_{\infty}\left(\frac{T}{\Delta}-k\right)},\frac{\delta}{2}\right).\end{split}$
@@
$P\left(\mathcal{L}_{g,T}(\hat{\eta}_{g,N,T})>\frac{\varepsilon}{6\left\|\frac{d(S^{-k}_{\Delta})_{*}\mu}{d\mu}\right\|_{\infty}\left(\frac{T}{\Delta}-k\right)}\right)\leq\frac{\delta}{2}$ (20)
 $\mbox{ for }N\geq N_{g}\left(\frac{\varepsilon}{6L_{\mathcal{H}_{G,N,T}}\left\|\frac{d(S^{-k}_{\Delta})_{*}\mu}{d\mu}\right\|_{\infty}\left(\frac{T}{\Delta}-k\right)},\frac{\delta}{2}\right).$ (21)
Using this along with (LABEL:eq:LipshitzEta_G) and (LABEL:eq:Est_g), we also see that the second term on the right hand side of (LABEL:eq:ErrorDecompFullState) is larger $\frac{\varepsilon}{2}$ with probability at most $\frac{\delta}{2}$ and for $N\geq N_{g}$. Therefore, for $N\geq\max\{N_{g},N_{G}\}$, we obtain the assertion of Theorem 5 from (LABEL:eq:ErrorDecompFullState), the union bound for probabilities and equations (20) and (21).
@@
Let $\Gamma$ be the quadratic lattice with lattice constant $\delta>0$, edge length $L$ and $\sqrt{k}=\frac{L}{\delta}$ lattice points on the edge. Eigenvectors of $A_{h}$ for $a=\text{const.}$ are those of the standard lattice Laplacian, which are well known. In fact, the trigonometric functions $(\sin(\omega\cdot\xi))_{\xi\in\Gamma}$, $(\cos(\omega\cdot\xi))_{\xi\in\Gamma}$ diagonalise $A_{h}$ where $\omega=(\omega_{1},\omega_{2})$ and $\omega_{j}\in\{2\frac{\pi j\delta}{L}:j\in\{0,\ldots,\sqrt{k}-1\}\}$.